how many gallons of pure alcohol should b mixed with 20 gallons of a 15% alchol solution to obtain a mixture that is 25% alcohol?
Found 2 solutions by stanbon, RAY100:
Answer by stanbon(75887)

how many gallons of pure alcohol should be mixed with 20 gallons of a 15% alcohol solution to obtain a mixture that is 25% alcohol?
--------------
alcohol + alcohol = alcohol
0.15*20 + x = 0.25(20+x)
Multiply thru by 100 to get:
15*20 + 100x = 25*20 + 25x
75x = 10*20
x = 2 2/3 gallons (amount of 100% alcohol needed for the mixture)
=====================================================================
Cheers,
Stan H.

Answer by RAY100(1637)

Let x = gallons of pure(100%) alcohol added
.
x(1.00) +20(.15) = (20 +x)(.25)
.
x+3=5+.25x,,,,,,,,,alcohol balance
.
,75 x = 2
.
x=2.67 gallons
.
.
check,,,2.67 +3 = 22.67 * .25 = 5.67,,,,,ok
.